A bridge that was 10.4 m long has been washed out by the rain several days ago. How fast must a car be going to successfully jump the stream, if we neglect air resistance? Although the road is level on both sides of the bridge, the road on the far side is 3.00 m lower than the road on this side.
Answer #1

time taken by car to cover the height

h = 0.5 gt^2

3 = 4.9* t^2

t = 0.78246 s

velocity needed

v = d/t

v = 10.4 / 0.78246

v = 13.291 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck